OpenStudy (anonymous):

Which of the following is the point-slope equation of the line with a slope= -5 and a point of (2, -1) ? WILL GIVE MEDAL AND FAN

OpenStudy (anonymous):

@Hero @jim_thompson5910 @k.rajabhishek @Chelsea04 @UnkleRhaukus

OpenStudy (anonymous):

@Kuoministers @saifoo.khan @skullpatrol

OpenStudy (anonymous):

D

OpenStudy (anonymous):

use the formula y-y1=m(x-x1)

OpenStudy (anonymous):

m as you know is -5

OpenStudy (anonymous):

then you substitute the point (2, -1) into the above equation, 2 is x1 -1 is y1

OpenStudy (anonymous):

so y-(-1)=-5(x-2)

OpenStudy (anonymous):

so then D. y+1=-5(x-2)

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

yeah :D you just substitute the numbers Thank you again!!

OpenStudy (anonymous):

no problem, here to help!